OpenStudy (anonymous):

A bucket is being filled at a rate of 5/6 gallon in 1/4 minute. What is the unit rate in gallons per minute? Both numbers are fractions!

OpenStudy (anonymous):

Are you saying the answer is both fractions or the fraction in the problem are both fractions?

OpenStudy (anonymous):

im saying 5/6 and 1/4 are both fractions

OpenStudy (anonymous):

Ok thanks for clearing up. Give me a second to explain

OpenStudy (anonymous):

Well, all we have to do is multiply 5/6 by 4, so we have \(\dfrac{5}{6}\times\dfrac{4}{1}\)

OpenStudy (anonymous):

How do you multiply that?

OpenStudy (anonymous):



OpenStudy (anonymous):

thank you sorry im late but its 3 and 1/3??